OpenStudy (anonymous):

Victor Malaba has a net income of $1,240 per month. If he spends $150 on food, $244 on a car payment, $300 on rent, and $50 on savings, what percent of his net income can he spend on other things? A. 40 percent B. 300 percent C. 43 percent D. 64 percent

OpenStudy (anonymous):

The total amount of money that he spends is: =150 + 244 + 300 + 50 = 744

OpenStudy (anonymous):

So the amount left with him is: = 1240 - 744 = 496

OpenStudy (anonymous):

as a percent, that would be = 496/1240 x 100

OpenStudy (anonymous):

*Hint* That's 0.4 * 100. Do the rest please @3dchick